# Strada-LLM: Graph LLM for traffic prediction

### v1

@@
Abstract
Traffic prediction is a vital component of intelligent transportation systems. By reasoning about traffic patterns in both the spatial and temporal dimensions, accurate and interpretable predictions can be provided. A considerable challenge in traffic prediction lies in handling the diverse data distributions caused by vastly different traffic conditions occurring at different locations. LLMs have been a dominant solution due to their remarkable capacity to adapt to new datasets with very few labeled data samples, i.e., few-shot adaptability. However, existing forecasting techniques mainly focus on extracting local graph information and forming a text-like prompt, leaving LLM-based traffic prediction an open problem. This work presents a probabilistic LLM for traffic forecasting with three highlights. We propose a graph-aware LLM for traffic prediction that considers proximal traffic information. Specifically, by considering the traffic of neighboring nodes as covariates, our model outperforms the corresponding time-series LLM. Furthermore, we adopt a lightweight approach for efficient domain adaptation when facing new data distributions in few-shot fashion. The comparative experiment demonstrates the proposed method outperforms the state-of-the-art LLM-based methods and the traditional GNN-based supervised approaches. Furthermore, Strada-LLM can be easily adapted to different LLM backbones without a noticeable performance drop.
Index Terms:
@@ Abstract
spatio-temporal transformers, graph neural networks, traffic prediction, LLM.
I Introduction
@@
With the development of Intelligent Transportation Systems, traffic forecasting has received increasing attention. It is a key component of advanced traffic management systems and is crucial in traffic planning, traffic management, and traffic control. Traffic forecasting involves analyzing traffic conditions on roads, including flow, speed, and density, mining traffic patterns, and predicting traffic trends [1]. This capability provides a scientific foundation for any traffic management department to anticipate and mitigate congestion, implement preemptive vehicle restrictions, and enable urban tourists to select safer and more efficient travel routes.
@@
Figure 1: An example of the spatio-temporal dependency between the traffic flows for different traffic states ①,②,③. Nodes (a), (b), and (c) are placed in a crowdedrea of Brussels. The corresponding traffic patterns between 8:00 am and 8:00 pm are illustrated. The short-term traffic flow similarity is altered by the significant impact of neighboring roads, resulting in varying flows.
@@
However, traffic forecasting is a challenging task due to complex spatial and temporal dependencies:
  1. 1.
@@
    Spatial Dependency: The variation in traffic volume is influenced by the topological structure of the urban road network. In particular, traffic conditions at upstream roads impact downstream roads through transfer effects, which refers to the impact that a change in one section of a traffic network (e.g., a closure of a road, a change in signal timing, or new infrastructure) has on other parts of the network. On the other hand, downstream roads affect upstream roads through feedback effects, which occur when the changes in traffic patterns resulting from transfer effects influence the original cause of the change, creating a loop of cause and effect.
  2. 2.
    Temporal Dependency: Traffic volume changes over time, exhibiting periodicity and trends, which are often influenced by factors such as holidays, working hours, and other social events.
@@
As illustrated in Figure 1, the strong influence among adjacent roads caused by the traffic flow alongside the edges changes the short-term state ① to state ② and state ③. Indeed, node (a) has a steady traffic count trend, and node (b) has short spikes in volume ②, so the spatio-temporal correlation caused state ③ in node (c) after a couple of hours. Furthermore, traffic data sources exhibit a variety of data distributions, posing a significant challenge for the domain in finding generalizable models.
@@
In recent years, the emergence of Large Language Models (LLM) have revolutionized various fields due to the generalization capabilities of extracting efficient features from different modalities [2, 3]. Research has been conducted in multiple domains, such as computer vision and natural language processing. Specifically, in  [4], a unified representation space is proposed, namely shared vocabulary, across varied data sources such that the model can learn domain-invariant features. This characteristic is particularly appealing for traffic prediction, where data distributions can vary significantly across different regions and times. In [5, 6, 7], the authors have tried to address this problem by casting the graph (and/or the corresponding time-series) to a prompt input. However, this branch of methods, in terms of predictability, is sub-optimal because traditional time series forecasting models use numerical inputs and outputs, whereas prompt-based methods transform these into text prompts, which may not be as effective for capturing the nuances of time series trends [8]. It will lead to a performance gap compared to classical GNNs [9, 10]. Moreover, the complexity and scalability of the prompt-based approaches are unexplored. For instance, in Spatio-Temporal Graphs(STG) with 1 million nodes, feeding the nodes to an LLM with limited context length is a challenging operation. Consequently, the prompting operation loses the global topological information. In [11], the structural information of the underlying graph is encoded as prompts for molecular analysis, but the effectiveness of this approach in urban contexts remains uncertain.
Parameter-Efficient Fine-Tuning (PEFT) has emerged as a favored method for integrating domain-specific expertise into Large Language Models (LLMs), particularly in the realms of language and vision processing [12]. This technique involves modifying a limited set of model parameters—either by targeting existing weights or introducing additional ones—while preserving the bulk of the model’s architecture. PEFT enables the retention of valuable pre-trained feature representations while concurrently tailoring the model to particular domains through targeted adjustments. This approach effectively strikes a balance between maintaining general knowledge and incorporating specialized insights. Among the popular PEFT strategies are LoRA [13] and prompt-tuning [14]. However, the optimal selection of an efficient method for an expert urban LLM, along with a comparative analysis of these techniques in this specific context, remains unexplored. Consequently, the question of how best to apply LLMs to graph-based traffic prediction using these methods is still open for investigation.
@@
To address these issues, we propose a new traffic forecasting method called Strada-LLM, a graph-aware LLM for traffic prediction. To the best of our knowledge, our model is the first attempt to propose a specialized parameter-efficient LLM model in traffic forecasting. Our contributions are five-fold:
  1. 1.
@@
    Strada-LLM is the first probabilistic LLM, which is compliant with classical LLMs and specializes in spatio-temporal traffic forecasting. Specifically, Strada-LLM has the capability to predict traffic, even when faced with a data distribution that differs from its training data, particularly in the few-shot setting.
  2. 2.
@@
    Strada-LLM is graph-aware, taking spatial dependencies into account by encoding the network graph implicitly, especially without any prompting.
  3. 3.
@@
    Strada-LLM adopts a lightweight approach to perform domain adaptation. Specifically, adopting low-ranked approaches showcases its effectiveness in adapting to new datasets.
  4. 4.
@@
    We provide a comprehensive evaluation and superior accuracy. We evaluate our approach using numerous real-world traffic datasets. Our results show a significant reduction in prediction error, ranging from approximately 5% to 18%, compared to baseline methods. This demonstrates the superiority of our graph-aware LLM model in traffic forecasting. We delve into the comparison between low-rank adaptation and prompt-tuning as two cogent approaches for LLM fine-tuning.
  5. 5.
    Strada-LLM is flexible and can be used with different known LLMs.
The paper is organized as follows. Section II examines pertinent studies related to traffic forecasting. Section III introduces the details of our method. In section IV, we evaluate the predictive performance of the Strada-LLM using real-world traffic datasets.
II Related Work
The related approaches can be categorized into two main groups: those specifically designed for traffic prediction and those developed for broader timeseries forecasting. The latter category includes methods for both univariate and multivariate time series analysis, which have been adapted to handle traffic-related variables.
II-A Statistical models
This class of models served as a fundamental pillar in traffic forecasting for many years, progressively evolving to tackle sophisticated forecasting challenges. Statistical models, such as ARIMA (Autoregressive Integrated Moving Average) [15] established the groundwork by utilizing auto-correlation to predict future outcomes. Subsequently, ETS (Error, Trend, Seasonality) models [16, 17] break down a time series into essential elements such as trends and seasonal variations, thus enabling more detailed forecasts. Theta models [18, 19] marked another notable progression by partially addressing non-linear behavior. Moreover,  [18, 20] often necessitates significant manual adjustment and specialized knowledge to choose suitable models and parameters for particular forecasting problems.
II-B Neural Traffic forecasting
Deep learning-based traffic forecasting is a rapidly developing research area [21]. Various architectures have been developed, including CNN-based [22, 10], RNN-based [23], and LSTM-based models[24, 25]. Inspired by the vanilla transformer mechanism, introduced in [26], PatchTST [27] highlights the importance of extracting local semantic information from time series data, which is often overlooked in previous models that use point-wise tokens, then re-process the output of a transformer for a point forecast or a probabilistic forecast. On the other hand, various other works propose alternative strategies to vanilla attention and build on the transformer architecture, leading to better models tailored for time series forecasting [28, 29, 30]. Specifically, models such as [31] propose spatio-temporal attention, others leverage trend-aware decomposition in the definition of the attention layer [32, 33, 34, 35, 3, 36]. On the other hand, there are copula-based transformer models in the literature [37], better at capturing volatilities, which are defined as the degree of variation in the time-series value over time. We selected transformer-based models due to their powerful ability to extract features and their effectiveness in few-shot learning scenarios. Specifically, we adopt Mistral [38] as our backbone model.
II-C LLMs
LLM models are an emerging paradigm within deep learning [39]. A key factor contributing to their popularity is the utilization of spatio-temporal attention [40], as highlighted by [31]. Many such models [41, 12] have showcased the power of robust feature extraction on a large scale. However, fine-tuning an LLM is time and energy-consuming, leading to the proposal of efficient methods [13, 42] with differing tuned parameters. LLMs have also been extended to scientific domains such as protein design [43, 44]. However, their application in urban traffic forecasting is still being investigated. Notably, several LLMs have been developed for traffic forecasting, including ST-LLM [45], which partially freezes the transformer block deteriorating the performance on a large scale. Moreover, Time-LLM [46], LLM4TS [47], and UniTime [48] cast the graph to textual prompt, that is, encode nodes, edges, attributes, etc., and forecast based on the prompts. Moreover, prompt-tuning emerged as a prominent technique for adapting large models [49]. FlashST [14] extends the prompt-tuning to spatio-temporal data. They adjust to unseen datasets by applying prompt-tuning in the context of spatio-temporal data. This category of models neglects the graph topology which hampers the prediction efficiency of the LLM [50]. Furthermore, data diversity has formed an additional challenge for LLMs in few-shot learning on novel datasets. [51]. The main goal of our work is to apply the LLM approach to traffic data forecasting and to investigate to what extent LLMs can adapt across a wide range of traffic datasets.
III Methodology
III-A Problem Definition
In this paper, we aim to predict traffic conditions at a specific timestamp $T{{}^{\prime}}$ based on historical traffic data over period $T$. Particularly, the traffic metric broadly refers to any traffic speed, flow, or density.
@@ III-A Problem Definition
Before further elaborating on our method, we define the following notations. In Strada-LLM, a road network is represented as a graph $G$, consisting of a set of nodes $V$ and a set of edges $E$. Each node presents a traffic measuring sensor, and an edge connects two nodes if they are geographically adjacent. That is, $G=(V,E,A)$ where $V$ is a set of $N$ nodes, $E$ is a set of edges and $A\in\mathbb{R}^{N\times N}$ is the corresponding adjacency matrix. We assume that the topology of the traffic graph $G$ is static. At each timestamp $t$, the graph $G$ is associated with a dynamic feature matrix $\boldsymbol{X}_{t}\in\mathbb{R}^{N\times F}$, where $F$ is the number of traffic metrics.
@@
In this paper, we formulate traffic forecasting as a discrete multi-variate point prediction problem. In particular, Strada-LLM takes the $T$ past observations $\boldsymbol{X}=\left[\boldsymbol{X}_{t_{1}},\boldsymbol{X}_{t_{2}},...,\boldsymbol{X}_{t_{T}}\right]\in\mathbb{R}^{T\times N\times F}$ from $G$ as input and predicts the traffic for horizon $T^{\prime}$, where $\small\boldsymbol{\hat{X}}=\left[\boldsymbol{\hat{X}}_{t_{T+1}},\boldsymbol{\hat{X}}_{t_{T+2}},...,\boldsymbol{\hat{X}}_{t_{T+T^{\prime}}}\right]$.
III-B Proposed Method
@@
Strada-LLM consists of a Hierarchical Feature Extractor (HFE), an LLM-based backbone, and a T-student distribution head. The distribution head includes three learnable parameters related to this distribution: degrees of freedom, mean, and scale, with suitable non-linearities to ensure the relevant parameters remain positive, as shown in Figure 2. In the HFE module, a sub-graph extractor and a global graph feature extractor hierarchically learn features while the LLM-based block maps the graph features into the latent spatio-temporal feature space to be learned by the distribution head. In the following sections, we will elaborate on each component.
III-B1 Hierarchical Feature Extractor
@@
  • Subgraph extractor: To enhance the LLM’s understanding of graph structures while accommodating its limited context length, we tokenize the spatio-temporal traffic signal regarding the existing subgraphs of the road network. Overall, the corresponding traffic signals for neighboring nodes will be aggregated together to be tokenized in the further steps. This block is responsible for extracting $k$-hop subgraphs from the input graph. For a node $v\in G$, the k-hop operator [52] is defined as:
    $\mathcal{N}_{k}(v)=\left\{u\in V|d(u,v)\leq k\right\},$ (1)
@@
    where $d(\cdot,\cdot)$ is the hop-based distance function between two nodes. By doing so, each node is equipped with a sub-graph of $G$, including local topological information. As a result, for each node $v$, we concatenate the features of $\mathcal{N}_{k}(v)$, leading to an $N\times T\times(M\times F)$ feature map, where $M$ denotes $|\mathcal{N}_{k}(v)|$. Similarly, Subgraph-1-WL [53] and MixHop [54] can also be used as the k-hop operator, but it is beyond the scope of this research.

### v3

@@
Abstract.
Traffic forecasting is pivotal for intelligent transportation systems, where accurate and interpretable predictions can significantly enhance operational efficiency and safety. A key challenge stems from the heterogeneity of traffic conditions across diverse locations, leading to highly varied traffic data distributions. Large language models (LLMs) show exceptional promise for few-shot learning in such dynamic and data-sparse scenarios. However, existing LLM-based solutions often rely on prompt-tuning, which can struggle to fully capture complex graph relationships and spatiotemporal dependencies—thereby limiting adaptability and interpretability in real-world traffic networks.
We address these gaps by introducing Strada-LLM, a novel multivariate probabilistic forecasting LLM that explicitly models both temporal and spatial traffic patterns. By incorporating proximal traffic information as covariates, Strada-LLM more effectively captures local variations and outperforms prompt-based existing LLMs. To further enhance adaptability, we propose a lightweight distribution-derived strategy for domain adaptation, enabling parameter-efficient model updates when encountering new data distributions or altered network topologies—even under few-shot constraints.
Empirical evaluations on spatio-temporal transportation datasets demonstrate that Strada-LLM consistently surpasses state-of-the-art LLM-driven and traditional GNN-based predictors. Specifically, it improves long-term forecasting by 17% in RMSE error and 16% more efficiency. Moreover, it maintains robust performance across different LLM backbones with minimal degradation, making it a versatile and powerful solution for real-world traffic prediction tasks.
@@ Abstract.
spatio-temporal transformers, graph neural networks, traffic prediction, LLM
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: August 03–07, 2025; Toronto, Canada; $c$s: Computing methodologiesccs: Machine learning approaches
1. Introduction
@@
With the development of Intelligent Transportation Systems, spatio-temporal traffic prediction has received increasing attention. It is a key component of advanced urban management systems and is crucial in urban planning, mobility management, and resource allocation. Spatio-temporal prediction involves analyzing urban conditions on various dimensions, including flow, speed, and density, mining their patterns, and predicting trends  (Li et al., 2018). This capability provides a scientific foundation for any urban management department to anticipate and mitigate congestion, implement preemptive restrictions, and enable urban residents to select safer and more efficient travel routes.
@@
Figure 1. An example of the spatio-temporal dependency between the traffic flows for different traffic states ①,②,③. Nodes (a), (b), and (c) are placed in a crowded area of Brussels. The corresponding traffic patterns between 8:$0$ am and 8:$0$ pm are illustrated for nodes (a), (b), and (c). The short-term traffic flow in (c) is altered by the significant impact of neighboring roads, resulting in varying flows.
@@
However, spatio-temporal prediction is a challenging task due to complex spatial and temporal dependencies:
@@
1) Spatial Dependency: The variation in urban patterns is influenced by the topological structure of the urban network. In particular, conditions at upstream locations impact downstream locations through transfer effects, which refers to the impact that a change in one section (e.g., a closure of an area, a change in service timing, or new infrastructure) has on other parts of the network. On the other hand, downstream locations affect upstream locations through feedback effects, which occur when the changes in patterns resulting from transfer effects influence the original cause of the change, creating a loop of cause and effect.
@@
2) Temporal Dependency: Urban conditions change over time, exhibiting periodicity and trends, which are often influenced by factors such as holidays, working hours, and other social events. As illustrated in Figure  1, the strong influence among adjacent locations caused by the flow alongside the edges changes the short-term states. Indeed, location (a) has a steady count trend, and location (b) has short spikes in volume, so the spatio-temporal correlation yielded a different state in location (c) after a couple of hours. Furthermore, urban data sources exhibit a variety of data distributions, posing a significant challenge for finding generalizable models.
@@
In recent years, the emergence of Large Language Models (LLM) has revolutionized various fields due to the generalization capabilities of extracting efficient features from different modalities  (Yuan et al., 2024). Research has been conducted in multiple domains, such as computer vision and natural language processing. Specifically, in  (Yuan et al., 2024), a unified representation space is proposed, namely shared vocabulary, across varied data sources such that the model can learn domain-invariant features. This characteristic is particularly appealing for urban prediction, where data distributions can vary significantly across different regions and times.
Previous works  (Li et al., 2024a, b) have tried to address this problem by casting the graph (and/or the corresponding time-series) to a prompt input. However, this branch of methods, in terms of predictability, is suboptimal because traditional time series forecasting models use numerical inputs and outputs, whereas prompt-based methods transform these into text prompts, which may not be as effective for capturing the characteristics and trends of time series  (Rasul et al., 2024). It leads to a performance gap compared to classical Graph Neural Networks  (Shao et al., 2022). Moreover, the complexity and scalability of the prompt-based approaches are unexplored. For instance, in Spatio-Temporal Graphs (STG) with 1 million nodes, feeding the nodes to an LLM with limited context length is a challenging operation. Consequently, the prompting operation loses the global topological information.
Effective LLM adaptation approach have a noticeable impact on their Zero-Shot performance. There are three prominent approaches to LLM adaptation(Alghisi et al., 2024). Namely, Prompt-Tuning(Qian et al., 2023), Adapter-Layer(Alghisi et al., 2024), Parameter-Efficient Fine-Tuning (PEFT)(Hu et al., 2021). The adapter-layer approach (Alghisi et al., 2024) tends to be more applicable in the multi-task setup for foundational models(Alghisi et al., 2024). In contrast, the effectiveness of Prompt-based approaches and PEFT remains a point of contention in the literature. Both methods have emerged as favored for integrating domain-specific expertise into Large Language Models (LLMs). However, in the realms of existing urban LLMs, UrbanGPT(Li et al., 2024a) and UniST(Yuan et al., 2024) have focused on Prompt-based approaches. Specifically, they include a learnable prompt network (PN) into the LLM. Essentially, the PN is fine-tuned in the adaptation stage. Prompt networks could impose extra complexity and neglect scalable features in the spatio-temporal target domain. Specifically, PN relies on the memory network(Sukhbaatar et al., 2015), which has been proven in graph-based spatio-temporal data to cause exponential error propagation(Barbero et al., 2024). On the other hand, PEFT(Hu et al., 2021) involves modifying a limited set of model parameters—either by targeting existing weights or introducing additional ones—while preserving the bulk of the model’s architecture. PEFT enables the retention of valuable pre-trained feature representations while concurrently tailoring the model to particular domains through targeted adjustments. However, the competence and scalability of PEFT in the LLM-based urban traffic forecasting domain remain unexplored.
@@
To address these issues, we propose a new forecasting method called Strada-LLM, which is a graph-aware LLM for urban traffic prediction. To the best of our knowledge, our model is the first to propose a specialized parameter-efficient LLM model in traffic forecasting. Our contributions are five-fold:
  1. (1)
@@
    Strada-LLM is a probabilistic LLM compliant with classical LLMs and specializes in spatio-temporal prediction. Specifically, Strada-LLM has the capability to predict urban traffic patterns, even when faced with a data distribution that differs from its training data, particularly in the few-shot setting.
  2. (2)
    To the best of our knowledge, Strada-LLM is the first non-prompt-based probabilistic LLM that has lightweight adaptation capability, effectively capturing both local and global spatial dependencies.
  3. (3)
@@
    Strada-LLM is graph-aware, taking spatial dependencies into account by encoding the network graph implicitly and without any prompting. This allows the model to maintain crucial topological information throughout the prediction process.
  4. (4)
@@
    Strada-LLM adopts a lightweight approach to perform domain adaptation. Specifically, adopting low-ranked approaches showcases its effectiveness in adapting to new datasets while maintaining computational efficiency.
  5. (5)
@@
    We provide a comprehensive evaluation and demonstrate superior accuracy compared to existing urban LLMs. We evaluate our approach using numerous real-world datasets. Our results show a reduction in prediction error, ranging from approximately 5% to 18%, compared to baseline methods. This demonstrates the superiority of our graph-aware LLM model in urban forecasting.
The remainder of this paper is organized as follows: Section 2 presents the preliminaries, Section 3 describes the methodology, Section 4 discusses the results, and Section 5 concludes.
2. Preliminaries
@@
Spatio-Temporal Forecasting. In this paper, we aim to predict traffic conditions at a specific timestamp $T{{}^{\prime}}$ based on historical traffic data over period $T$. Particularly, the traffic metric broadly refers to traffic speed, flow, or density. Before further elaborating on our method, we define the following notations. In Strada-LLM, a road network is represented as a graph $G$, consisting of a set of nodes $V$ and a set of edges $E$. Each node presents a traffic measuring sensor, and an edge connects two nodes if they are geographically adjacent. That is, $G=(V,E,A)$ where $V$ is a set of $N$ nodes, $E$ is a set of edges and $A\in\mathbb{R}^{N\times N}$ is the corresponding adjacency matrix. We assume that the topology of the traffic graph $G$ is static. At each timestamp $t$, the graph $G$ is associated with a dynamic feature matrix $\boldsymbol{X}_{t}\in\mathbb{R}^{N\times F}$, where $F$ is the number of traffic metrics.
@@
In this paper, we formulate traffic forecasting as a discrete multi-variate point prediction problem. In particular, Strada-LLM takes the $T$ past observations $\boldsymbol{X}=\left[\boldsymbol{X}_{t_{1}},\boldsymbol{X}_{t_{2}},...,\boldsymbol{X}_{t_{T}}\right]\in\mathbb{R}^{T\times N\times F}$ from $G$ as input and predicts the traffic for horizon $T^{\prime}$, where:
$\boldsymbol{\hat{X}}=\left[\boldsymbol{\hat{X}}_{t_{T+1}},\boldsymbol{\hat{X}}_{t_{T+2}},...,\boldsymbol{\hat{X}}_{t_{T+T^{\prime}}}\right]$}.
3. Methodology
@@
Strada-LLM consists of a Hierarchical Feature Extractor (HFE), an LLM-based backbone, and a T-student distribution head, as shown in Figure 2. The distribution head includes three learnable parameters related to the output: degrees of freedom, mean, and scale, to ensure the relevant parameters remain positive. In the HFE module, a sub-graph extractor and a global graph feature extractor hierarchically learn features while the LLM-based block maps the graph features into the latent spatio-temporal feature space to be learned by the distribution head. In the following sections, we will elaborate on each component.
3.1. Hierarchical Feature Extractor
3.1.1. $k$-hop Sub-Graph Extractor:
@@
To enhance the LLM’s understanding of graph structures while accommodating its limited context length, we tokenize the spatio-temporal traffic signal by taking into account the existing subgraphs of the road network. Overall, the corresponding traffic signals for neighboring nodes will be aggregated together to be tokenized in the further steps. This block is responsible for extracting $k$-hop subgraphs from the input graph. For a node $v\in G$, the k-hop operator (Nikolentzos et al., 2019) is defined as:
(1) $\mathcal{N}_{k}(v)=\left\{u\in V|d(u,v)\leq k\right\},$
@@
where $d(\cdot,\cdot)$ is the hop-based distance function between two nodes. By doing so, each node is equipped with a sub-graph of $G$, including local topological information. As a result, for each node $v$, we concatenate the features of $\mathcal{N}_{k}(v)$, leading to an $N\times T\times(M\times F)$ feature map, where $M$ denotes $|\mathcal{N}_{k}(v)|$, cardinality of neighbors. Similarly, Subgraph-1-WL (Zhou et al., 2024) and MixHop (Abu-El-Haija et al., 2019) can also be used as the k-hop operator, but it is beyond the scope of this research.
3.1.2. Global Graph Embedding